Last year, X Company sold 69,200 units of its only product for $17.00 each. Total costs were as follows: Cost of goods sold Variable $424,196 Fixed 131,480 Selling and administrative Variable $76,120 Fixed 96,188 At the end of the year, a company offered to buy 4,090 units of the product but only for $11.00 each. X Company had the capacity to produce the additional units, and even though there would have been no additional selling and administrative costs, it rejected the offer.

If X Company had accepted the special order, firm profits would have increased by?

X Company's production manager thought that the special order units might have required additional direct material and direct labor costs per unit of $0.75 and $0.29, respectively, and the rental of special equipment for $1,000. The combined effect of these changes would have been to reduce the special order profit by?

X Company's marketing manager felt that in order for regular sales to continue at 69,200 units this year, the company would have had to reduce the regular selling price this year to $16.53. The result of this price reduction would have been to reduce X Company's profits this year by?
Answer #1

Solution 1:

Variable cost of goods sold per unit = $424,196 / 69200 = $6.13 per unit

If X company accept the special order then profit will increase by = ($11 - $6.13) * 4090 = $19,918

Solution 2:

Reduction in special order profit by = ($0.75 + $0.29) * 4090 + $1,000 = $5,254

Solution 3:

Price reduction would reduce X company profit by = (69200*$0.47) = $32,524